OpenStudy (anonymous):

What is the slope of the line represented by the equation y = 8

OpenStudy (anonymous):

y=mx+c m is your slope. Your C here is zero so what's your m?

OpenStudy (anonymous):

y=mx+c m is your slope. Your C here is zero so what's your m?

OpenStudy (anonymous):

y=mx+c m is your slope. Your C here is zero so what's your m?

OpenStudy (lgbasallote):

y = 8 is a horizontal line horizonta lines ALWAYS have 0 slope

Parth (parthkohli):

\( \color{Black}{\Rightarrow y = mx + b }\) We see that here is only one possiblility, i.e: \( \color{Black}{\Rightarrow y = 0x + 8 }\) \(m\) is the slope, right? So m = 0

OpenStudy (anonymous):

Slope of a y = ? is always 0 since it's a horizontal line. Slope of a x = ? is always undefined because a vertical line does not have slope.

OpenStudy (anonymous):

Your C is 8. apologies.

Parth (parthkohli):

As 0x = 0, we may say that this is in the form y = 0x + 8

OpenStudy (anonymous):

Your C is 8. apologies.

OpenStudy (anonymous):

Your C is 8. apologies.

Parth (parthkohli):

@beeqay :)